OpenStudy (anonymous):

Log(3x+4) = 1 + log (x-6). Solve for x.

OpenStudy (anonymous):

I moved both logs to one side so: Log(3x+4) - log (x-6) = 1 and then I did that thing \[\frac{ \log(3x+4) }{ \log(x-6) }=1\]

OpenStudy (anonymous):

But I don't know what to do now...

OpenStudy (tkhunny):

That's good that you don't know what to do next, because what you did, did not work. "That thing" should be this: \(\log(3x+4) - \log(x-6) = \log\left(\dfrac{3x+4}{x-6}\right)\). Now, what's next?

OpenStudy (anonymous):

That's what I have written down on paper I just wrote it wrong here, I swear. Thanks for pointing it out, though.

OpenStudy (anonymous):

I know what I have to do, but I don't know HOW to do it. Like, if this was raised to an exponent I'd get the square root. I don't know what to do for logs...

OpenStudy (anonymous):

My terminology is off, I know. I'm sorry.

OpenStudy (anonymous):

Third time I've attempted this and I keep getting it wrong.

OpenStudy (anonymous):

Mhmm?

OpenStudy (wolf1728):

log [(3x+4)/(x-6)] = 1

OpenStudy (anonymous):

ask yourself when does the log of something =1 then solve the equation for that number

OpenStudy (anonymous):

But I'm supposed to solve for x.

OpenStudy (wolf1728):

or to put it another way, what is the anti-log of 1?

OpenStudy (anonymous):

\[10 = \frac{ 3x+4 }{ x-6 }\]

OpenStudy (wolf1728):

Very good Atlas!

OpenStudy (anonymous):

okay let w= all the crap in the parenthesis log(w)=1 where does log of w=1 once you figure that out let that number=(3x+4)/(x-6)

OpenStudy (anonymous):

and solve for x

OpenStudy (anonymous):

I knew it was something simple. I hate myself. lol

OpenStudy (anonymous):

Just forgot how logs worked.

OpenStudy (anonymous):

Thanks everyone.

OpenStudy (anonymous):

okay your welcome.

OpenStudy (wolf1728):

u r welcome :-)